# How to set AdamW’s weight decay as you scale model and dataset size

### v1

@@ Abstract
We show that weights learned by AdamW can be understood as an exponential moving average (EMA) of recent updates. This gives critical insights for how to set the weight decay in AdamW, and how the weight decay should scale with model and dataset size. In particular, the key hyperparameter for an exponential moving average is the EMA timescale. Intuitively, the EMA timescale can be understood as the number of recent iterations the EMA averages over. Given a fixed learning rate, there is a one-to-one mapping from the EMA timescale to the usual weight decay hyperparameter. Thus, choosing an EMA timescale implicitly sets the weight decay. Importantly, there are natural guidelines for sensible values for the EMA timescale: we need to average over all datapoints, so the EMA timescale should not be (much) smaller than 1 epoch, and we need to forget early updates, so the EMA timescale should not be (much) bigger than the total number of training epochs. In our experiments, we find that optimal EMA timescales are consistent with these guidelines, as are the hyperparameters chosen in recent large-scale LLM pretraining runs (e.g. Llama 1+2 and Stable LM). Critically, these guidelines suggest that the optimal EMA timescale should not change (much) as we scale the model and dataset. That implies that as the dataset size increases, the optimal weight decay should fall. Moreover, as the model size increases, the optimal weight decay should also increase (if we follow the muP recommendation for scaling the learning rate).
@@ 1 Introduction
A common machine learning workflow is to prototype by training many smaller models, then at the end to do one final training run, with the largest possible model on the largest possible dataset. This workflow is used in many settings, from small scale student projects to the largest LLM training runs. However, for this workflow to be effective, we need to understand how to transfer optimal hyperparameters from smaller-scale prototyping runs to the final, largest model and dataset. In this paper, we focus on transferring the weight decay hyperparameter in AdamW (Loshchilov & Hutter, 2018), as AdamW is used in many settings including the largest LLM pretraining runs (e.g. Zhang et al., 2022; Touvron et al., 2023a, b; Tow et al., 2023).
The most obvious approach for changing the hyperparameters as we scale the model size (but not the dataset size) is µP (Yang et al., 2022). The key recommendation from µP is to scale the learning rate, $\eta$, as $1/\text{fan\_in}$. However, the theory behind µP analyses only how the learning rate affects behaviour close to initialization as we scale model size; this theory says nothing about the effect of regularization as the optimizer approaches convergence. As such, µP does not provide the whole answer to hyperparameter transfer in AdamW, given the need to additionally transfer weight decay hyperparameters, $\lambda$, across model and dataset sizes.
@@
To understand how to set the AdamW weight decay and how to transfer it across model and dataset sizes, we show that AdamW can be understood as an exponential moving average (EMA). Of course, both Adam and AdamW use EMAs to estimate the average gradient, $m_{t}$, and average squared gradient, $v_{t}$. That is not what we are talking about. Instead, we observe that in algorithms with decoupled weight decay (i.e. AdamW but not Adam), the weights themselves are an EMA of recent updates (see Sec. 3.1 for details).
The key hyperparameter in an EMA is the EMA timescale, which intuitively describes the number of previous iterations that the EMA averages over; this timescale is $\tau_{\text{iter}}=1/(\eta\lambda)$ (Sec. 3.1). Thus, setting the learning rate, $\eta$, and timescale, $\tau_{\text{iter}}$, implies a value for the weight decay, $\lambda=1/(\eta\tau_{\text{iter}})$. As such, given a fixed learning rate, $\eta$, we can implicitly set the weight decay, $\lambda$, by setting the EMA timescale, $\tau_{\text{iter}}$. We argue that it is easier to reason about and tune the EMA timescale, $\tau_{\text{iter}}$, because there is a natural range of values for $\tau_{\text{iter}}$. Specifically, if we measure the timescale in epochs, i.e. $\tau_{\text{epoch}}=\tau_{\text{iter}}/M$, where $M$ denotes the number of iterations per epoch, then $\tau_{\text{epoch}}$ should not be far smaller than 1, otherwise the EMA would not average over all the datapoints. At the same time, the $\tau_{\text{epoch}}$ should not be far larger than the total number of training epochs, otherwise the algorithm would be unable to forget contributions to the EMA from very early on in the training process.
We perform a number of experiments showing this natural range for the optimal $\tau_{\text{epoch}}$ seems to capture the actual optimal $\tau_{\text{epoch}}$ (Sec. 4.1). Moreover, we note that this natural range of values has been used for successful large-scale LLM training runs such as Llama 1+2 (Touvron et al., 2023a, b) and Stable LM 3B (Tow et al., 2023).
Next, we notice that this natural range of values for $\tau_{\text{epoch}}$ does not seem to change with model and dataset size. That has important implications for how to scale the weight decay with the model and dataset size. Specifically, as the dataset size increases, we find that the optimal weight decay falls (Sec. 4.2). Moreover, if we follow the usual µP recommendation for the learning rate as we increase the model sizes, we find that the optimal weight decay should increase as the model size increases (Sec. 4.3). We validate both of these predictions experimentally.
@@ 2 Background
Yang et al. (2022) considered how to transfer the learning rates in SGD and Adam across model sizes. They considered two key desiderata. First, the initial random weights, $W^{0}\in\mathbb{R}^{D\times\text{fan\_in}}$, times an input, $x\in\mathbb{R}^{\text{fan\_in}}$, should not grow or shrink as the model changes size. We could write this requirement as, $W^{0}x\sim 1$, where we use $\sim$ to indicate “scales with”. This gives the usual $W^{0}\sim 1/\sqrt{\text{fan\_in}}$ scaling of the standard deviation of the initial random weights. At the same time, they required that the change in the outputs, caused by the first weight update, $\Delta Wx$, shouldn’t grow or shrink as the model changes size; i.e. $\Delta Wx\sim 1$. Yang et al. (2022) show that this requirement implies that the learning rate should scale as $1/\text{fan\_in}$.
The distinction between the $1/\sqrt{\text{fan\_in}}$ scaling of the initial weights vs. the $1/\text{fan\_in}$ scaling of the learning rate might be a bit puzzling. The intuitive reason for this distinction is given in Yang et al. (2022) Appendix J. In short, for the initial weights,
$y_{i}=\sum_{j=1}^{\text{fan\_in}}W^{0}_{ij}x_{j}$ (1)
@@
$\Delta y_{i}=\sum_{j=1}^{\text{fan\_in}}\Delta W_{ij}x_{j},$ (2)
thus, $\Delta y_{i}\sim\text{fan\_in}\times\Delta W_{ij}$. To ensure that $\Delta y_{i}\sim 1$, we therefore need $\Delta W_{ij}\sim 1/\text{fan\_in}$. As the Adam updates, $\Delta W_{ij}$ scale with the learning rate, $\eta$, this implies $\eta$ must also scale with $1/\text{fan\_in}$.
@@
$\Delta W_{ij}$ $=\eta\frac{\hat{m}_{ij}}{\sqrt{\hat{v}_{ij}}}$ (3)
where we have neglected the small $\epsilon$. Here, $\hat{m}_{ij}$ is an EMA estimate of the gradient $g_{ij}$, while $\hat{v}_{ij}$ is an EMA estimate of the expected squared gradient $g_{ij}^{2}$. Thus, $\hat{m}_{ij}\sim g_{ij}$ and $\hat{v}_{ij}\sim g_{ij}^{2}$. That implies $\hat{m}_{ij}/\sqrt{\hat{v}_{ij}}\sim 1$, so looking back at Eq. 3, we have $\Delta W_{ij}\sim\eta$. Hence, to get $\Delta W_{ij}\sim 1/\text{fan\_in}$, we need $\eta\sim 1/\text{fan\_in}$, which is the origin of the $1/\text{fan\_in}$ scaling of the learning rates in µP.
Importantly, as Yang et al. (2022) considers the size of the first update relative to the random initial weights, it does not give guidance about how to change the weight decay with the model size, as the weight decay only becomes relevant after many learning steps.
@@ 3.1 AdamW as an EMA
$w_{t}$ $=(1-\eta\lambda)w_{t-1}-\eta\frac{\hat{m}_{t}}{\sqrt{\hat{v}_{t}}+\epsilon}$ (4)
@@
where $w_{t}$ is a neural network weight, $\eta$ is the learning rate, $\lambda$ is the weight decay, $\epsilon$ is a small constant, $\hat{m}_{t}$ is a bias-correct EMA estimate of the expected gradient, and $\hat{v}_{t}$ is a bias-correct EMA estimate of the expected squared gradient. Notice that here we adopt the parameterization used by major optimization libraries such as torch.optim [^1] https://pytorch.org/docs/stable/generated/torch.optim.AdamW.html and optax [^2] https://optax.readthedocs.io/en/latest/api/optimizers.html#optax.adamw, where the decay for $w_{t-1}$ is controlled by the product $\eta\lambda$. This parameterization is different from the form suggested in the original AdamW paper (Loshchilov & Hutter, 2018), which we will discuss in detail in the related work section.
@@
Now we will show that these weight updates (Eq. 4) can be understood as an EMA. Recall that a generic exponential moving average estimate, $\text{ema}_{t}$ can be written as,
$\operatorname{ema}_{t}$ $=(1-\nicefrac{{1}}{{\tau_{\text{iter}}}})\operatorname{ema}_{t-1}+\nicefrac{{1}}{{\tau_{\text{iter}}}}\;q_{t}.$ (5)
@@
Here, $\operatorname{ema}_{t}$ forms an exponential moving average estimate of $q_{t}$, and the EMA timescale is $\tau_{\text{iter}}$. Specifically, if we take the canonical form for an EMA (Eq. 5) and set:
@@
$\nicefrac{{1}}{{\tau_{\text{iter}}}}$ $=\eta\lambda$ $\operatorname{ema}_{t}$ $=w_{t}$ $q_{t}$ $=-\frac{1}{\lambda}\frac{\hat{m}_{t}}{\sqrt{\hat{v}_{t}}+\epsilon}$ (6)
@@
then we recover the AdamW updates (Eq. 4). Of course, AdamW uses EMAs to compute $\hat{m}_{t}$ and $\hat{v}_{t}$. Our key insight was that additionally, the overall AdamW updates for $w_{t}$ can themselves be understood as EMAs.
Importantly, there are only two hyperparameters under the EMA view: the EMA timescale, $\tau_{\text{iter}}=1/(\eta\lambda)$ and the initialization, $\operatorname{ema}_{0}$. The connection between AdamW and EMA thus suggests that AdamW also should have only two hyperparameters, $\tau_{\text{iter}}$, and something about the initialization. In Appendix A we show just this. Specifically, for scale-invariant models, the full trajectory of an AdamW optimizer depends only on $\tau_{\text{iter}}$, and the relative size of the parameter initialization against the learning rates.
3.2 Implication of the EMA interpretation
The previous section argued that AdamW is an EMA, and the key hyperparameter in an EMA is the timescale, $\tau_{\text{iter}}$. Here, we argue that the EMA viewpoint suggests a range of sensible values for the EMA timescale. We can rewrite AdamW’s EMA as a weighted average of all updates (see Appendix B for a derivation of this form).
@@ 3.2 Implication of the EMA interpretation
$w_{t}$ $=\nicefrac{{1}}{{\tau_{\text{iter}}}}\sum_{t^{\prime}=1}^{t}e^{-(t-t^{\prime})/\tau_{\text{iter}}}\left(-\frac{1}{\lambda}\frac{\hat{m}_{t^{\prime}}}{\sqrt{\hat{v}_{t^{\prime}}}+\epsilon}\right)$ (7)
@@
where we make a particular choice of the (unnormalized) EMA weights, $e^{-(t-t^{\prime})/\tau_{\text{iter}}}$ . These weights highlight that recent updates (i.e. $t^{\prime}$ close to $t$) contribute more to the weighted average. Specifically, for recent updates where $t-t^{\prime}$ is far smaller than $\tau_{\text{iter}}$, the weights, $e^{-(t-t^{\prime})/\tau_{\text{iter}}}$, are around $1$. At the same time, updates from further into the past contribute less: when $t-t^{\prime}$ is far larger than $\tau_{\text{iter}}$, then the weights, $e^{-(t-t^{\prime})/\tau_{\text{iter}}}$, decay to zero. Critically, this implies that the EMA averages over roughly the last $\tau_{\text{iter}}$ minibatch updates.
However, it is still difficult to understand the implications of averaging over the previous $\tau_{\text{iter}}$ iterations. Instead, we propose to work with the timescale measured in terms of epochs,
$\tau_{\text{epoch}}=\tau_{\text{iter}}/M.$ (8)
where $M$ is the total number of minibatches in the dataset or equivalently iterations in an epoch.
Now, $\tau_{\text{epoch}}$ has a natural interpretation. Specifically, if $\tau_{\text{epoch}}$ is around $1$, then the AdamW EMA averages over updates in the last epoch. Alternatively, if $\tau_{\text{epoch}}$ is far smaller than $1$, then the AdamW EMA averages over only a small fraction of the training samples, which is likely to be very suboptimal. Likewise, if $\tau_{\text{epoch}}$ is too big (compared to the total number of epochs in the full training run), AdamW will be unable to forget the random initialization and/or very early gradients, which again could lead to suboptimal performance.
4 Results
In this section, we first demonstrate that the optimal value of $\tau_{\text{epoch}}$ falls within a natural range. We then show that as dataset sizes vary, the optimal $\tau_{\text{epoch}}$ remains within a narrow range for the same model. Finally, we illustrate that across different model sizes, the optimal timescale also remains consistent within a close range.

### v2

@@ Abstract
The scaling of the optimal AdamW weight decay hyperparameter with model and dataset size is critical as we seek to build larger models, but is poorly understood. We show that weights learned by AdamW can be understood as an exponential moving average (EMA) of recent updates. This gives critical insights for how to set the weight decay in AdamW, and how the weight decay should scale with model and dataset size. In particular, the key hyperparameter for an exponential moving average is the EMA timescale. Intuitively, the EMA timescale can be understood as the number of recent iterations the EMA averages over. We find that the optimal timescale, measured in epochs, is roughly constant as we change model and dataset size. Moreover, given a learning rate, there is a one-to-one mapping from the EMA timescale to the weight decay hyperparameter. Thus, if the optimal EMA timescale is constant, that implies that as the dataset size increases, the optimal weight decay should fall and as the model size increases, the optimal weight decay should increase (if we follow the muP recommendation for scaling the learning rate). We validate these scaling rules on ResNet-18 and Vision Transformers trained on CIFAR-10 and ImageNet, and on NanoGPT pre-training on OpenWebText. Finally, we found that as training progresses, muP’s learning rate scaling breaks down for AdamW unless weight decay is scaled appropriately.
Machine Learning, ICML
@@ 1 Introduction
A common machine learning workflow is to prototype by training many smaller models, then at the end do one final training run, with the largest possible model on the largest possible dataset. This workflow is used in many settings, from small-scale student projects to the largest LLM training runs. However, for this workflow to be effective, we need to understand how to transfer optimal hyperparameters from smaller-scale prototyping runs to the final, largest model and dataset. This is very well studied for the optimal learning rate as you scale up width and depth (e.g. Yang et al., 2022; Bordelon et al., 2024; Noci et al., 2024; Everett et al., 2024). However, the largest LLM pretraining runs (e.g. Zhang et al., 2022; Touvron et al., 2023a, b; Tow et al., 2023) use an optimizer with a weight decay term such as AdamW (Loshchilov & Hutter, 2018), and it is not understood how to scale the AdamW weight decay hyperparameter with model sizes, and how to scale hyperparameters as we scale dataset size. Here, we fix this gap, showing how the optimal weight decay transfers across model width and dataset size.
@@
To understand how to transfer weight decay across model and dataset sizes, we argue that AdamW should be understood as an Exponential Moving Average (EMA). Of course, both Adam and AdamW use EMAs to estimate the average gradient, $m_{t}$, and average squared gradient, $v_{t}$. That is not what we are talking about. Instead, we observe that in algorithms with decoupled weight decay (i.e. AdamW but not Adam), the weights themselves are an EMA of recent updates (see Sec. 3.1). The key hyperparameter in an EMA is the EMA timescale, which intuitively describes the number of previous iterations that the EMA averages over, which is given by $\tau_{\text{iter}}=1/(\eta\lambda)$ (Sec. 3.1). In fact, we prove and empirically verify that under a $\eta$-dependent initialization and a fully-scale invariant network, AdamW’s optimization trajectory is controlled solely by the timescale $\tau_{\text{iter}}$ for all combinations of learning rate $\eta$ and weight decay coefficient $\lambda$ (Sec. 3.2).
Perhaps more importantly, the EMA perspective tells us how to set the weight decay parameter. Specifically, we can take into account the dataset and batch sizes to measure the timescale in epochs, i.e. $\tau_{\text{epoch}}=\tau_{\text{iter}}/M$, where $M=N/\mathrm{B}$ denotes the number of iterations per epoch in a training dataset of $N$ samples under a batch size of $\mathrm{B}$, which measures the fraction of the dataset that contributes to the current weights through the EMA. We empirically observed that the optimal $\tau_{\text{epoch}}$ is stable across model and dataset sizes, which, consequently provides a guideline for scaling weight decay. In particular, as the dataset size $N$ increases, we find that the optimal weight decay falls (Sec. 4.1). Moreover, if we follow the usual µP recommendation to decrease the learning rate $\eta$ as the model width increases, we find that the optimal weight decay should increase as the model size increases (Sec. 4.2). We validate both of these predictions in ResNet, vision transformer trained on CIFAR-10 and ImageNet, and NanoGPT pre-training on OpenWebText.
In summary, our paper’s contributions are:
  • The weights generated by AdamW can be written as an EMA of recent weight updates.
  • The optimal EMA timescale changes little as we scale the model and dataset size.
[…]
  • The optimal weight decay increases with model width.
  • When using AdamW with fixed weight decay, µP learning rate scaling breaks down, but proper weight decay scaling restores its effectiveness
@@ 2 Background
[…]
$y_{i}={\textstyle\sum}_{j=1}^{\text{fan\_in}}W^{0}_{ij}x_{j}$ (1)
@@
$\Delta y_{i}={\textstyle\sum}_{j=1}^{\text{fan\_in}}\Delta W_{ij}x_{j},$ (2)
[…]
@@
$\Delta W_{ij}$ $=\eta\hat{m}_{ij}/\sqrt{\hat{v}_{ij}}$ (3)
[…]
Importantly, as Yang et al. (2022) considers the size of the first updates relative to the random initial weights, it does not give guidance about how to change the weight decay with the model size, as the weight decay only becomes relevant after many learning steps.
@@ 3.1 AdamW as an EMA
$w_{t}$ $=(1-\eta_{t}\lambda)w_{t-1}-\eta_{t}\frac{\hat{m}_{t}}{\sqrt{\hat{v}_{t}}+\epsilon}$ (4)
@@
where $w_{t}$ is a neural network weight, $\eta_{t}$ is the learning rate (which potentially varies over time due to scheduling), $\lambda$ is the weight decay, $\epsilon$ is a small constant, $\hat{m}_{t}$ is a bias-corrected EMA estimate of the expected gradient, and $\hat{v}_{t}$ is a bias-corrected EMA estimate of the expected squared gradient. Notice that here we adopt the parameterization used by major optimization libraries such as torch.optim [^1] https://pytorch.org/docs/stable/generated/torch.optim.AdamW.html, where the decay for $w_{t-1}$ is controlled by the product $\eta_{t}\lambda$. This parameterization is different from the form suggested in the original AdamW paper (Loshchilov & Hutter, 2018), which we will discuss in detail in the related work section. Further, we follow the usual convention, established by Loshchilov & Hutter (2018) of scheduling the learning rate but not the weight decay.
@@
Now we will show that these weight updates (Eq. 4) can be understood as an EMA [^2] For an introduction to EMAs, see en.wikipedia.org/wiki/Exponential_smoothing. Recall that a generic exponential moving average estimate, $\text{ema}_{t}$ can be written as,
$\operatorname{ema}_{t}$ $=(1-\nicefrac{{1}}{{\tau_{\text{iter};t}}})\operatorname{ema}_{t-1}+\nicefrac{{1}}{{\tau_{\text{iter};t}}}\;q_{t}.$ (5)
@@
Here, $\operatorname{ema}_{t}$ forms an exponential moving average estimate of $q_{t}$, where the potentially time-varying EMA timescale is $\tau_{\text{iter};t}$. We take this EMA (Eq. 5) and set:
@@
$\nicefrac{{1}}{{\tau_{\text{iter};t}}}$ $=\eta_{t}\lambda$ $\operatorname{ema}_{t}$ $=w_{t}$ $q_{t}$ $=-\frac{1}{\lambda}\frac{\hat{m}_{t}}{\sqrt{\hat{v}_{t}}+\epsilon}$ (6)
@@
then we recover the AdamW updates (Eq. 4). Of course, AdamW uses EMAs to compute $\hat{m}_{t}$ and $\hat{v}_{t}$. Our key insight was that additionally, the overall AdamW updates for $w_{t}$ can themselves be understood as EMAs. Intuitively, the AdamW EMA computes the average of $q_{t}$, i.e. the noisy minibatch update scaled by the weight decay.
To build intuition, we consider a fixed timescale, $\tau_{\text{iter}}$, and write the EMA as a weighted average of all minibatch updates (derivation in Appendix B),
@@
$w_{t}$ $=\nicefrac{{1}}{{\tau_{\text{iter}}}}\sum_{t^{\prime}=1}^{t}e^{-(t-t^{\prime})/\tau_{\text{iter}}}\left(-\frac{1}{\lambda}\frac{\hat{m}_{t^{\prime}}}{\sqrt{\hat{v}_{t^{\prime}}}+\epsilon}\right).$ (7)
@@
Thus, the unnormalized weights are $e^{-(t-t^{\prime})/\tau_{\text{iter}}}$. These weights highlight that recent updates (i.e. $t^{\prime}$ close to $t$) contribute more to the weighted average. Specifically, for recent updates where $t-t^{\prime}$ is far smaller than $\tau_{\text{iter}}$, the weights, $e^{-(t-t^{\prime})/\tau_{\text{iter}}}$, are around $1$. At the same time, updates from further into the past contribute less: when $t-t^{\prime}$ is far larger than $\tau_{\text{iter}}$, then the weights, $e^{-(t-t^{\prime})/\tau_{\text{iter}}}$, decay to zero. Thus, intuitively, the EMA averages over roughly the last $\tau_{\text{iter}}$ minibatch updates.
3.2 Formalizing intuition from the EMA
In an EMA, perhaps the key parameter is the timescale, $\eta_{t}\lambda$. Additionally, the initial value is also important; in AdamW, this corresponds to the size of the first update relative to the scale of the initialization, $\eta_{0}/\sigma$. Indeed, Theorem 1shows that, under a scale-invariant network, if these two quantities are fixed, the whole optimization trajectory is determined.
Theorem 1.
Consider two AdamW optimizers with different learning rates, weight decays, initialization scales and epsilons, ($\eta_{t},\lambda,\sigma,\epsilon$ vs. $\eta_{t}^{\prime},\lambda^{\prime},\sigma^{\prime},\epsilon^{\prime}$). Take $w_{t}$ to be the parameters learned by the first optimizer after the $t$th optimization step, and $w_{t}^{\prime}$ to be the parameters learned by the second optimizer. These optimizers are initialized by scaling the same random init,
$w_{0}$ $=\sigma{\xi}$ $w_{0}^{\prime}$ $=\sigma^{\prime}{\xi},$ (8)
where $\xi$ is random noise (e.g. IID Gaussian). Consider a scale-invariant network, in the sense that multiplying the weights, $w$, by an arbitrary positive constant, $01/c>$ 0 gives the same output for all inputs, $x$,
$\operatorname{net}(x;w)$ $=\operatorname{net}(x;\tfrac{1}{c}w).$ (9)
We use the same trajectory of EMA timescales, $\nicefrac{{1}}{{\tau_{\text{iter};t}}}=\eta_{t}\lambda=\eta_{t}^{\prime}\lambda^{\prime}$, the same ratio of the first step, relative to the initial parameter scale, $\tfrac{\eta_{0}}{\sigma}=\tfrac{\eta_{0}^{\prime}}{\sigma^{\prime}}$. Now, the entire trajectory of weights for each optimizer is the same (up to a scalar multiplier,) $w_{t}^{\prime}=\tfrac{1}{c}w_{t}$ where $c=\tfrac{\sigma}{\sigma^{\prime}}=\tfrac{\eta}{\eta^{\prime}}=\tfrac{\lambda^{\prime}}{\lambda}=\tfrac{\epsilon^{\prime}}{\epsilon}.$
Thus, the network outputs are the same at all points along the trajectory, $\operatorname{net}(x;w_{t})=\operatorname{net}(x;w_{t}^{\prime})$. For the proof, see Appendix A.1, and for empirical confirmation, see Appendix A.3. Thus, for scale-invariant networks, only two parameters are relevant to the trajectory: the EMA timescale (and its schedule), $\tau_{\text{iter};t}$, and the scale of the initial learning rate relative to the weight init. We would usually expect the influence of the init to diminish over time, leaving $\tau_{\text{iter};t}$ as the key hyperparameter.
4 Scaling weight decay across model and dataset sizes
Since timescale is the key hyperparameter in an EMA, we hypothesize that the optimal timescale should be transferrable across model and dataset sizes. In this section, we empirically verify this hypothesis and consider the implications for the optimal weight decay as we increase the training dataset size (Eq. 11) and model size (Eq. 16).